OpenStudy (ayeeeitskateee):

Can somebody help me solve a 30-60-90 triangle?

Mehek (mehek14):

what's the question

OpenStudy (ayeeeitskateee):

Bottom left. I need to find x and y.

OpenStudy (ayeeeitskateee):

And top left I also need help with. I really don't know how to do them at all. lol.

OpenStudy (goformit100):

User Trigonometrical ratios concepts, and, pytharoras Theorem

OpenStudy (ayeeeitskateee):

Still not sure how to do it, lol. ^

OpenStudy (ayeeeitskateee):

@iGreen Do you know how to do this? lol.

OpenStudy (anonymous):

Solve the following for x and y.:\[\left\{x^2=y^2+\left(8 \sqrt{3}\right)^2,x=2 y\right\} \]

OpenStudy (ayeeeitskateee):

Oh lord, that looks confusing.

OpenStudy (anonymous):

Replace x^2 in the first equation with 4 y^2 and then solve the result for y.

OpenStudy (anonymous):

With a 30 deg triangle, y is 1/2 of x.